### v1

# Cloud2BIM: An open-source automatic pipeline for efficient conversion of large-scale point clouds into IFC format

Abstract
Building Information Modeling (BIM) is an essential component in the sustainable reconstruction and revitalization of ageing structures. However, model creation usually relies on laborious manual transformation of the unstructured point cloud data provided by laser scans or photogrammetry. This paper presents Cloud2BIM, an open-source software tool designed to automate the conversion of point clouds into BIM models compliant with the Industry Foundation Classes (IFC) standard. Cloud2BIM integrates advanced algorithms for wall and slab segmentation, opening detection, and room zoning based on real wall surfaces, resulting in a comprehensive and fully automated workflow. Unlike existing tools, it avoids computationally- and calibration-intensive techniques such as RANSAC, supports non-orthogonal geometries, and provides unprecedented processing speed—achieving results up to seven times faster than fastest competing solutions. Systematic validation using benchmark datasets confirms that Cloud2BIM is an easy-to-use, efficient, and scalable solution for generating accurate BIM models, capable of converting extensive point cloud datasets for entire buildings into IFC format with minimal user input.
keywords:
3D Reconstruction , Automation , BIM , Python , Classification
journal: Journal
1 Introduction
Over the last decade, the construction industry has experienced a profound digital transformation, marking a new architectural and engineering design epoch. This period, commonly referred to as Construction 4.0 (Sawhney et al., 2020), is characterized by a pivotal shift toward the utilization of advanced digital technologies aimed at enhancing the efficiency, quality, safety, and sustainability of construction projects (Statsenko et al., 2023).
Amid these advancements, there is a growing focus on sustainability, especially for buildings nearing the end of their lifecycle. This focus has spurred construction companies to develop and adopt innovative methods that significantly reduce the environmental impacts of building and demolition activities. In Europe, construction and demolition account for approximately 36% of total waste generation, highlighting a considerable environmental challenge (Eurostat, 2020).
In response to the demand for sustainable construction practices, the industry is increasingly retrofitting and modernizing existing buildings. These initiatives typically involve upgrading insulation, installing energy-efficient fixtures, and integrating renewable energy systems to enhance building performance and sustainability (Dakwale et al., 2011; Grecchi, 2022). Furthermore, adopting deconstruction and recycling practices provides a practical alternative by reducing waste and decreasing the dependence on new, raw materials (Ge et al., 2017). Although deconstruction can be slower and more costly than traditional demolition (Bertino et al., 2021), it is crucial for transitioning the construction industry towards a circular economy.
However, the effective implementation of these sustainable strategies is often hampered by the absence or obsolescence of building documentation. Building models, essential for documentation, can be accurately created using laser scanning or photogrammetry. Both methods generate detailed three-dimensional coordinates and potentially color information of points on the surfaces of structures. The density of these point clouds varies depending on the technology used and the distance from which the scans are taken. Photogrammetry employs conventional RGB images to construct point clouds (Aber et al., 2019), whereas laser scanning utilizes laser beams to capture comprehensive 3D data, forming detailed point clouds (Valero et al., 2022).
Moreover, as Mirzaei et al. (2023) report, these point clouds are invaluable for extracting precise geometry information about individual building members, aiding in their dimensional quality control during construction or renovation projects. Despite their high accuracy and visually precise representations, point clouds are inherently unstructured and do not support database operations such as listing individual elements, nor do they contain contextual information. In contrast, entity-based representations like building information models (BIM) incorporate detailed metadata and relational data within each component, enabling thorough analysis and manipulation. The capability to transform point clouds into precise, digital, entity-based models facilitates the planning of retrofits or reconstructions with enhanced accuracy, supporting conservation and sustainable development goals (Akbarieh et al., 2020).
Converting point cloud data into usable BIM, or scan-to-BIM process, is often a meticulous and labor-intensive task that demands significant expertise and advanced software (Klinc et al., 2021; Abreu et al., 2023). This procedure is predominantly managed by skilled BIM modelers manually integrating point cloud data with parametric objects. While this manual approach is practical, it is also highly time-consuming and results in considerable costs. This manual labor requirement presents a challenge, particularly for small investors, as the high costs of manually created 3D models and the subsequent generation of BIM are often prohibitive despite the potential value of such models. Recent developments have introduced both semi-automatic (Dore and Murphy, 2014) and fully automatic (Kwadjo et al., 2019; Romero-Jarén and Arranz, 2021; Wang et al., 2015; Martens and Blankenbach, 2023; Mahmoud et al., 2024) solutions to address this issue.
The available scan-to-BIM solutions exploit various algorithms for segmenting and classifying building elements. For segmenting slabs, analyzing point density along the vertical axis using height histograms, as explored in a few studies (Oesau et al., 2014; Turner and Zakhor, 2014; Romero-Jarén and Arranz, 2021; Martens and Blankenbach, 2023), became a standard method. This approach is simple yet sensitive to noise, which can lead to inaccuracies. As an alternative, Nikoohemat et al. (2018) introduced an algorithm that detects horizontal planes by analyzing trajectories from mobile laser scanners, offering effective results but requiring specific scanning technology.
Wall classification often employs the random sample consensus (RANSAC) algorithm, which identifies sets of points that align with a plane, frequently utilized for detecting interior walls (Schnabel et al., 2007; Ochmann et al., 2016; Murali et al., 2017; Ochmann et al., 2019). Another common strategy simplifies scenes to a 2D plan view, analyzing wall positions under the assumption that walls are vertical, which simplifies processing but restricts the geometrical possibilities (Oesau et al., 2014; Bormann et al., 2016; Macher et al., 2017; Jung et al., 2018).
Results from wall geometry analysis can be categorized into two main approaches: (i) room-based, which focuses on the dimensions of spaces and rooms (Wang et al., 2017; Yang et al., 2019), and (ii) wall-based or volumetric, which concentrates on wall geometry and the 3D model consisting of wall objects (Ochmann et al., 2019; Nikoohemat et al., 2018; Bassier and Vergauwen, 2020; Martens and Blankenbach, 2023; Jung et al., 2018). The room-based approach typically identifies only the surfaces of rooms, whereas the wall-based approach involves identifying opposing surfaces that form walls and solving their topology and connections.
Despite the plethora of semi-automated and automated scan-to-BIM solutions available, a fully open-source, universally applicable tool independent of commercial software is still lacking in the current literature. While some initiatives have made source code available, such as the notable work of Bassier and Vergauwen (2020), these examples are the exception rather than the rule. The lack of widely available code hinders the ability of the construction industry to adopt and adapt new technologies and methodologies effectively. In addition, many existing solutions rely on proprietary software to function, such as the Dynamo add-on [^1] https://dynamobim.org/ with Revit software by Chen (2022) and Mahmoud et al. (2024), or depend on the Revit API as documented by Mehranfar et al. (2024b).
This paper introduces Cloud2BIM, a novel Python-based open-source tool that fills the gap in the current scan-to-BIM landscape. It provides a fully automated and flexible conversion process, free from dependencies on commercial software. Unlike existing methods, Cloud2BIM handles complex architectural geometries, such as non-orthogonal wall configurations, and efficiently processes large data sets with minimal user input. A distinctive feature of Cloud2BIM is its ability to export data directly in IFC format, ensuring interoperability across different BIM software platforms. It also includes advanced zoning capabilities that more accurately define space characteristics using actual wall surfaces. By simplifying the conversion process and offering it as an open-source solution, Cloud2BIM makes advanced BIM technology available to a broader audience, promoting sustainable construction practices through precise and rapid transformation of raw scan data into usable BIM outputs.
2 Testing datasets
Three distinct datasets were chosen to evaluate the proposed solutions and ensure an unbiased assessment of algorithm performance. The datasets comprised two own scans conducted by our team and one synthetic benchmark dataset. A Trimble X7 scanner was employed for the scans. A professional surveyor meticulously created these scans by establishing a geodetic network, stabilizing field points, determining coordinates with GNSS technology, and refining the network through terrestrial surveying. The resulting scans were of high quality and required no accuracy modifications, although they were thinned to reduce computational demands.
The process of data dilution was executed using the open-source CloudCompare software, where the point clouds were diluted to a prescribed minimum distance to the nearest point ($d_{\mathrm{min}}$). In ’spatial’ mode, the user specifies a minimum distance between points. CloudCompare then selects points from the original cloud, ensuring no point in the resulting cloud is closer to another than the specified distance. As this value increases, fewer points are retained. The underlying algorithm operates as follows: Points are processed in an order determined by their spatial organization within an octree structure. The initial point is marked as ’to be kept’. For each subsequent point, the algorithm checks for neighbors within the specified radius. If no neighboring points are marked as ’to be kept’, this point is also marked. Ultimately, only points marked as ’to be kept’ are included in the final output cloud. The impact of this dilution is illustrated in Figure 1.
Such an approach to dilution of point clouds proved effective for buildings with relatively simple geometric details. However, for buildings of historical significance that feature intricate decorations, the dilution strategy proposed by Štroner et al. (2022) is more suitable. This method adjusts the dilution factor based on the local curvature, helping to preserve critical features while reducing the risk of feature loss, often referred to as ’feature washing.’ Caution is advised with this approach as it tends to remove points from planar surfaces, which could lead to insufficient density for effective point cloud segmentation. Although avoiding dilution prevents data loss, it dramatically increases computational demands and may not always be practical.
2.1 ETH Zürich synth3
The synthetic dataset, provided by the Multimedia Laboratory at the University of Zurich222www.ifi.uzh.ch/en/vmml/research/datasets.html, offers a complex challenge for testing the construction of volumetric walls under the Atlanta-world assumption (Table 1). This assumption permits the depiction of scenes using vertical and horizontal planes in 3D, without the constraint of vertical planes (walls) being orthogonal to each other (Pintore et al., 2020). The dataset encompasses 20 interior walls, features 13 door openings, and is designed for algorithms that reconstruct interior walls connecting at various angles. Notably, the external surfaces of the object were not created by the dataset authors, simulating purely indoor scanning. This dataset has been pivotal in research focused on indoor wall detection, contributing significantly to other studies (Ochmann et al., 2019; Shi et al., 2019; Bassier and Vergauwen, 2020).
Dataset name Visualization $n_{\mathrm{points}}$ Size $d_{\mathrm{min}}$
ETH Zürich synth3 [Uncaptioned image] 6,905,684 468 kB 0.01
Table 1: ETH Zürich synth3 dataset specification.
2.2 Kladno station
Point cloud data from Kladno railway station (Table 2), captured using the terrestrial laser scanner, was employed to demonstrate the wall segmentation capabilities of the developed software on a real-world dataset. For this purpose, one floor of the building was utilized. Using an actual point cloud from this building provides the benefit of including real-world irregularities and imperfections that the program must navigate, thereby enhancing the robustness and applicability of the segmentation results.
Dataset name Visualization $n_{\mathrm{points}}$ Size $d_{\mathrm{min}}$
Kladno station [Uncaptioned image] 2,551,721 190 kB 0.025
Table 2: Kladno station dataset specification.
2.3 Prague hotel
The point cloud dataset representing a former hotel in Prague (Table 3) encompasses an eleven-story building section. This dataset was chosen to demonstrate the software’s ability to manage large datasets and to highlight the precision and efficiency improvements afforded by density-based height segmentation. The laser scanning was conducted from inside the building using a terrestrial laser scanner, resulting in the building envelope being largely absent, except for areas visible through window openings.
Dataset name Visualization $n_{\mathrm{points}}$ Size $d_{\mathrm{min}}$
Prague hotel [Uncaptioned image] 40,152,849 2.54 GB 0.025
Table 3: Prague hotel dataset specification.
Figure 1: Histogram of minimum distances to the nearest neighbor before (left) and after (right) dilution; ETH Zürich synth3 dataset (top), Kladno station dataset (middle), and Prague hotel dataset (bottom).
3 Methods
This section outlines the algorithms used to obtain the geometric representation of building structures through Cloud2BIM, our Python-based software that adheres to open-source principles. A comprehensive list of the program dependencies and packages utilized during development is provided in Table 4. The entire codebase is accessible on our GitHub repository. Figure 2 presents a diagram illustrating the program workflow.
Figure 2: Flowchart illustrating the process of generating a 3D model of a building from laser scans, detailing each step from data acquisition to model completion.
Cloud2BIM supports two types of file formats for input: .e57 and .xyz. These formats generally provide information about the position of individual points $(x,y,z)$. While the .E57 format is designed to include colors (r, g, b) and intensities as standard attributes, the .XYZ format may include colors and intensities, but this is not standardized and depends on the specific implementation. For accurate classification within Cloud2BIM, only the position data of the points are utilized.
Table 4: List of Python and associated packages used in the development of the Cloud2BIM software, including their versions and types of licenses.
Software/Library Version License
Python 3.10 PSF license
Numpy 1.26.2 BSD license
Matplotlib 3.8.1 PSF license
Pye57 0.4.1 MIT license
Pandas 2.1.3 BSD 3-Clause license
Open3D 0.17.0 MIT license
Tqdm 4.66.1 MIT license
Scikit-image 0.22.0 BSD 3-Clause license
Opencv-python 4.8.1.78 MIT license
Scipy 1.11.3 BSD 3-Clause license
Ifcopenshell 0.7.0.231018 GNU lesser general public license
3.1 Identification of slabs
Cloud2BIM adopts a widely used method that analyzes point density within a histogram to identify horizontal planes in point clouds (Bassier and Vergauwen, 2020), illustrated in Figure 3. This approach assumes that the floor slab is horizontal relative to the $x$-$y$ plane, has uniform thickness, and lacks local height variations like steps. A candidate for a horizontal surface is identified within a height range that contains at least 50% of the maximum number of points found in any single plane. This threshold is adjustable to account for uneven point distribution often seen in real-world scans, where fewer points may be recorded at floor level due to obstructions like furniture or scanner positioning.
Figure 3: Point density histogram (left) showing point counts by height ($z$-coordinate) with a 50% maximum threshold marked by a dashed blue line. The side view of the point cloud (right) displays point distribution in the $x$-$y$ plane, identifying horizontal surface candidates (orange) above the threshold and other points (blue) below it.
The floor-plan shape of the slab is then determined based on the surface index:
  1. 1.
    The first slab, typically representing either a ground-level or a partially scanned slab, is assumed to have only its upper surface and requires a manual thickness setting.
  2. 2.
    If an even number of surfaces remains, it implies that both the ceiling slabs and possibly the roof have been scanned from both surfaces. Here, the surfaces are merged to create a floor plan point cloud for robust analysis, assuming that the slab mirrors the ceiling shape of the floor below (Figure 4).
  3. 3.
    An odd number of surfaces suggests that the last surface is unpaired, potentially indicating an unscanned or non-flat roof, which also requires a manual thickness setting.
Figure 4: Candidate point cloud subsets for merging representing the floor surface (left) with lower density due to obstructions during scanning and ceiling surface (right) used to create a floor plan when combined.
To process the point cloud, the function create_hull_from_histogram projects the 3D points onto a 2D plane, discarding the $z$-coordinate, and creates a 2D histogram (Figure 5) that assists in forming a binary mask. Morphological operations—dilation and erosion—are then applied to refine the mask contours, maintaining object shape while smoothing irregularities. The mask is adjusted to ensure precise contour alignment with the grid.
Figure 5: A 2D histogram of cloud point density in the identified floor plane (ETH Zürich synth3 dataset).
Using the algorithm by Suzuki and Abe (1985) implemented through OpenCV’s findContours() function, the most prominent contour is extracted, scaled, and converted into a Polygon object that delineates the floor plan. To simplify this contour, the Douglas-Peucker algorithm (Douglas and Peucker, 1973) is applied, reducing the number of points and smoothing the shape, as shown in Figure 6. This process effectively represents the polygon of a slab with fewer, more significant points. The complete program workflow for slab segmentation is depicted in Figure 7.
Figure 6: A close-up view of the ceiling slab edge, with the original edge indicated in blue and the smoothed edge, generated using the Douglas Peucker algorithm in red.
Figure 7: Flowchart diagram illustrating the function used to detect slabs in an unorganized point cloud.
3.2 Segmentation into storeys
Once the horizontal structures are identified, the point cloud is segmented into individual storeys. This segmentation facilitates the spatial organization of the building. The segmentation process is handled by the function split_pointcloud_to_storeys, which requires two primary inputs: points_xyz for the 3D coordinates of the point cloud, and slabs (Figure 8). Each slab is represented by a dictionary detailing the polygon of the slab, its $x$- and $y$-coordinates, its lower $z$-coordinate, and its thickness. The function processes each slab sequentially, extracting points between the base of one slab and the top of the slab below it while maintaining by a default setting 10 cm clearance from both the ceiling and the floor. The output is a list of point cloud subsets, where each subset corresponds to a distinct storey within the building.
Figure 8: Flowchart diagram illustrating the process of segmenting a point cloud into individual storeys, highlighting the sequence of operations from slab identification to storey-specific point cloud extraction.
3.3 Identification of walls
Two primary strategies prevail for reconstructing interior walls from point cloud data: room-based and wall-based. The room-based approach simplifies the segmentation process by focusing on reconstructing the geometry of individual rooms, which involves segmenting the point cloud into distinct spaces and determining the optimal shape for each. This approach, however, struggles with flexible parametric wall geometry and unenclosed spaces. On the other hand, the wall-based approach clusters point cloud data around each wall, aiming to reconstruct topologically consistent structures (Bassier and Vergauwen, 2020).
The Cloud2BIM’s Identify_walls function, described using the flowchart in Figure 9, detects and defines wall geometry within a point cloud. It isolates individual wall segments within each floor’s horizontal section, set by default to 90–100%
Figure 9: Flowchart diagram illustrating the algorithms implemented in the Identify_walls function for segmentation of walls.
This 2D point cloud then undergoes histogram analysis to create a 2D histogram, where each bin’s value reflects the point density in that spatial region. Converting this histogram into a binary image based on a predefined threshold isolates regions with high point densities (likely indicative of walls). This binary image (Figure 10) serves as the basis for further wall detection and geometry definition. It undergoes processing to extract contours using the findContours function from the OpenCV library (Bradski, 2000), explicitly targeting the outermost boundaries of wall segments. Each detected contour is then approximated into line segments through the Douglas-Peucker algorithm. These segments are converted from pixel-based coordinates to real-world coordinates.
Figure 10: Binary mask highlighting high-density regions indicative of wall positions, derived from a 2D histogram of the point cloud (top) and contours delineating wall segments in the 2D floor plan, derived from the binary mask segmentation (bottom); ETH Zürich synth3 dataset.
Following segmentation, the wall segments undergo further analysis to join into longer segments based on their mutual angles, distances, and alignment. To form a volumetric representation, opposite wall surfaces are paired if they are parallel, within a maximum distance, and have sufficient overlap. A virtual surface is generated to complete the pairing if an outer surface is missing.
The final steps involve calculating the wall axis and adjusting connections to ensure correct topology. The algorithm identifies the longest segment as a reference, calculates its direction vector, and uses it to estimate the central position of the wall axis. Endpoints of wall segments are adjusted for intersections within the threshold of half the maximum wall thickness to ensure walls connect seamlessly.
The function outputs each wall’s start and end points along with its thickness. The wall height is assumed to be the full height of the floor, as this method does not allow for the detection of shorter walls.
3.4 Identification of zones
Zones are entities related to the internal areas of individual rooms, typically bounded by floor slabs, ceilings, and surrounding walls. Zones facilitate the extraction of area and volumetric parameters for individual rooms and enable the assignment of various attributes related to their function, cladding materials, fire-safety parameters, acoustics, and other relevant aspects.
The detection of zones is typically approached through three primary methods (Chen et al., 2024): (i) slice-based modeling, (ii) mesh simplification modeling, and (iii) floorplan-based modeling. Once the exact positions and geometries of walls are identified, Cloud2BIM circumvents these computationally intensive methods. Instead, the identify_zones function (Figure LABEL:fig:zones_flochart) analyzes wall geometry by first processing the axes of walls to ensure connectivity and proper alignment. Wall segments are split at intersections of these axes and extended to connect with adjacent axes within a distance specified by the snapping_distance parameter. Subsequently, segments parallel to the wall axes, representing the surfaces of the walls, are generated at a distance equal to half the width of the individual walls. These parallel segments are then extended to intersect fully, forming the boundaries of zones. The function ultimately creates closed polygons defined by the wall axes; each polygon serves as a candidate zone. The complete process of zone identification is illustrated in Figure 12.
Figure 11: Flowchart diagram illustrating the identify_zones function.
Figure 12: Wall axes after snapping to the nearest neighbours are disjoined at intersections (top images), polygons created based on the identified wall axes (middle), and final boundaries of identified zones bounded by the corresponding wall surfaces (bottom), demonstrated on the ETH Zürich synth3 (left) and Kladno station (right) datasets.
3.5 Identification of openings
Several methods for detecting openings have been developed over the years. Techniques for identifying door openings have typically been categorized into histogram-based approaches (Mahmoud et al., 2024; Chen, 2022) and those that utilize scanner trajectory data (Nikoohemat et al., 2018). Each method, however, comes with specific limitations. Density-based analysis can be sensitive to obstructions like furniture, affecting the histogram’s accuracy, while trajectory-based methods depend heavily on particular scanning technology. Due to these constraints, our approach utilizes a density analysis from histograms, as illustrated in the flowchart in Figure 13, where the function identify_openings is outlined.
Figure 13: Flowchart diagram illustrating the identify_openings function.
Several heuristic rules enhance the histogram-based method in Cloud2BIM to pinpoint door and window openings accurately. These rules consider the sill position to differentiate between windows and doors, the aspect ratio of the opening, and the maximum and minimum widths of the opening. However, this approach restricts the detected shapes to rectangular openings.
The analysis begins by positioning the wall at the coordinate origin and rotating it into the $x$-$z$ plane. The floor plan position of the opening, primarily along the $x$-axis, is first determined. The histogram is then analyzed based on a threshold defined as a percentage of the tenth maximum value recorded to identify the start and end of gaps or low-density areas. The tenth maximum was selected based on observations of the histograms and found appropriate to avoid extreme maxima that seldom occur in the histogram; in some walls there were about 1–3 extreme maxima so that the tenth maximum provides a safe reserve, but the number of selected maximum can be changed. When matching openings are detected, this procedure is repeated for the $z$-axis histogram. Figure 14 presents a side view of the wall, displaying histograms for the $x$ and $z$ axes with marked door openings in red. The effectiveness of this analysis is improved by the prior dilution of the point cloud to the minimum distance, as discussed in Section 2. A more uniform point distribution minimizes the risk of misjudging the thresholds for empty space in the histograms, thereby improving the accuracy of opening detection.
Figure 14: Histogram-based wall opening detection, featuring a side view of the wall (top left), point density histograms along the x-axis (bottom left) and z-axis at the opening location (top right), with dashed lines marking the threshold for opening detection.
3.6 Creation of IFC output file
The industry foundation classes (IFC) file format is primarily designed to facilitate data transfer among various participants in the modeling process, allowing the sharing of the building information model and its associated data. Utilizing IFC as the final format to transfer data from point clouds to BIM is a well-established approach, as evidenced by similar implementations by Martens and Blankenbach (2023), which included the reconstruction of slabs, floors, and walls. There are numerous scan-to-BIM methodologies discussed in the literature, such as those by Adán et al. (2018) and Truong-Hong and Lindenbergh (2022); however, these approaches typically do not support direct export to IFC or any standardized BIM format. The capability to export directly to IFC, as implemented in our Cloud2BIM tool, facilitates seamless integration with a wide range of BIM software platforms. This compatibility significantly enhances the utility of the generated models for architectural planning, structural analysis, and lifecycle management, ensuring that stakeholders can leverage the full spectrum of BIM functionalities for efficient project execution and maintenance.
Our output of Cloud2BIM is generated in the IFC 4 ADD2 TC1 format using the IfcOpenShell library for Python (IfcOpenShell contributors, 2023). This process includes specifying all necessary attributes for each IFC class. We chose the DesignTransferView_V1.0 Model View Definition to ensure that our models are compatible with and can be further edited in BIM software platforms such as ArchiCAD [^3] https://graphisoft.com/solutions/archicad, Revit [^4] https://www.autodesk.com/products/revit, and Allplan [^5] https://www.allplan.com.
The structural hierarchy of IFC elements mirrors standard project organization practices as each IFC file must define precisely one instance of IfcProject or another IfcContext, such as IfcProjectLibrary, which serves as the anchor for all other data within the file. IfcProject includes essential attributes such as unit definitions, project name, and project phase, while other IFC classes like IfcSite, IfcBuilding, and IfcBuildingStorey help in assigning objects to their precise locations. In terms of geometry, the IfcSlab class is used to represent ceiling slabs, where the slab’s perimeter is defined by an IfcArbitraryClosedProfileDef, typically a polyline and extruded along a specified axis using the IfcExtrudedAreaSolid entity. Walls are similarly modeled using the IfcWall class, with geometry defined by start and end points via the IfcCartesianPoint class, and extruded using the IfcExtrudedAreaSolid. Openings in the walls are handled by the IfcOpening class, linked to the parent wall through the IfcRelVoidsElement relationship, with their positions defined by the IfcAxis2Placement3D and modeled using the swept solid technique. Zones are created using the same techniques as slabs using the Ifcspace class. They are based on IfcArbitraryClosedProfileDef, which defines the area of the zone, and IfcExtrudedAreaSolid, which defines the height of the zone extrusion. Graphical representation of these classes is provided in Figure 15.
Figure 15: The most important IFC classes used for geometry representation.
3.7 Inputs and calibration parameters
For accurate model configuration, Table 5 provides a comprehensive list of input parameters, encompassing details about the point cloud, the scanned building, and additional project information that cannot be derived directly from the laser scan. These parameters ensure that the software is precisely tailored to the specifics of each project. Table 6 details the calibration values essential for the effective classification and segmentation of building elements from the point cloud data, optimizing the accuracy of the generated model.
Table 5: Input parameters for the Cloud2BIM software, detailing settings related to the point cloud, building elements, and project details.
Parameter Unit Related object/entity
pc_resolution m point cloud
bfs_thickness m slab
tfs_thickness m slab
min_wall_length m wall
min_wall_thickness m wall
max_wall_thickness m wall
exterior_walls_thickness m exterior wall
snapping_distance m wall axis, zone
material_for_objects material
ifc_site_latitude degree site
ifc_site_longitude degree site
ifc_site_elevation m site
ifc_project_name project
ifc_project_long_name project
ifc_project_version project
ifc_author_name author
ifc_author_surname author
ifc_author_organization author
ifc_building_name building
ifc_building_type building
ifc_building_phase building
Table 6: Calibration parameters for the segmentation and classification algorithms, including default values and descriptions for each setting to optimize model accuracy; the default values were set based on testing the algorithms based on different datasets, but in the near future the software will be equipped with sliders and graphical interface to allow convenient setting of these parameters.
Parameter Unit Default value Description
dilution_factor 10 Controls the skipping of rows in the XYZ point cloud file.
grid_coefficient px/mm 5 Sets the grid size for computational operations in binary image processing.
z_step m 0.05 Defines the bin size for the horizontal histogram in slab surface detection.
max_n_points_array 0.5 Multiplier for maximum point density to identify $z$-coordinates for slabs.
dilation_meters m 1 Specifies the dilation distance for morphological operations on slab masks.
erosion_meters m 1 Specifies the erosion distance for morphological operations on slab masks.
smoothing_factor m 0.0005 Determines the smoothing level of slab polygon contours using the Douglas-Peucker algorithm.
safety_margin m 0.1 Margin (overlap) used when splitting the point cloud into storeys.
z_section_boundaries [0.9, 1.0] Sets the height percentage thresholds for segmenting point clouds for wall detection.
threshold 0.01 Sets the value threshold for creating a positive mask from 2D histograms.
square(5) px 5 Kernel size and shape for adjusting the wall detection mask.
epsilon m/px 0.02 Controls the approximation of line segments.
angle_tolerance degree 3 Tolerance level for merging collinear segments during wall reconstruction.
max10 10 Sets the threshold for determining if a histogram bin represents an opening.
4 Results and discussion
4.1 Comparative analysis and benchmarking of algorithms
To contextualize our software’s capabilities within the field, Table 7 provides a comparative analysis of recent developments by other researchers tackling similar challenges in the point-cloud-to-BIM conversion. This table highlights essential criteria for classifying, segmenting, and reconstructing complete building point clouds, including the automation of multi-floor segmentation, the creation of volumetric elements like slabs, walls, and zones, and the handling of complex architectural features such as oblique wall axes and content elements like furniture. The criteria also cover the capabilities of opening detection, stairs modeling, and providing output in IFC file format, alongside the availability of the complete open-source software code. Comparative accuracies are not presented due to the lack of a uniform benchmark dataset and undisclosed accuracy metrics in many referenced studies.
Table 7: A comparative analysis of the functionality of solutions reported in recent articles focused on the point cloud to model conversions. Both recent and older studies are included to provide a comprehensive overview, reflecting the ongoing and historical contributions to this specialized field.
Multi-story
Slabs
Oblique wall axes
Wall surfaces
Volumetric walls
Volumetric spaces
Openings
Stairs
Content elements
IFC output
Open-source
Wang et al. Wang et al. (2015)
Macher et al. Macher et al. (2017)
Jung et al. Jung et al. (2018)
Shi et al. Shi et al. (2019)
Ochmann et al. Ochmann et al. (2019)
Bassier and Vergauwen Bassier and Vergauwen (2020)
Romero-Jarén and Arranz Romero-Jarén and Arranz (2021)
Gen and Gentes Chen and Gentes (2021)
Martens and Blankenbach Martens and Blankenbach (2023)
Mahmoud et al. Mahmoud et al. (2024)
Mehranfar et al. Mehranfar et al. (2024a)
Our solution (Cloud2BIM)
In reviewing the contributions of others, our discussion emphasizes the distinction in methodologies, scope, and outcomes. We highlight how different authors have approached the segmentation of point clouds and the geometric reconstruction of architectural elements, pointing out the strengths and limitations of their approaches compared to ours. For instance, Bassier and Vergauwen (2020) focuses solely on walls, creating geometry in Revit through a Rhino plug-in and then exporting to IFC, including zones. Our method supports direct IFC export from Python, reducing dependence on commercial software. In addition, their approach, like some others, lacks the ability to detect and model doors and windows, often resulting in a lower level of detail. This study though, unlike the other, allows
The work of Mahmoud et al. (2024) incorporates advanced deep learning techniques for semantic segmentation, followed by RANSAC for line extraction to identify architectural features. While it provides detailed segmentation of walls, floors, and ceilings, the geometry export is limited to an Excel format, which is then used by a Dynamo add-in for Revit, constraining the output to specific commercial software. Similarly, the work by Mehranfar et al. (2024b) relies on the Revit API for 3D geometry and employs the PointTransformer network with DBScan clustering. However, this approach is sensitive to auxiliary parameter settings and does not enhance functionality beyond our method. Furthermore, it only supports single-floor point clouds, limiting its utility for multi-storey buildings.
Martens and Blankenbach (2023) presented VOX2BIM+, a robust approach by analyzing voxels to improve noise resistance and accuracy in the volumetric reconstruction of spaces and structural elements. Unlike our approach, which utilizes direct point cloud processing, their method incorporates several advanced techniques, including region growing and morphological skeletons for wall detection. However, their code remains proprietary, and their method does not support window or door geometry extraction.
Similarly, Wang et al. (2015) and Ochmann et al. (2019) have made significant contributions to the field, focusing on energy simulation models to detailed wall segmentation. These studies typically rely on simplified models or specific algorithms like RANSAC, which may only capture some architectural details such as doors and windows, a capability our software aims to address comprehensively. While sufficient for their goals, the detected zones that are clustered according to their purpose and not published to IFC cannot be further edited and exported to other software.
4.2 Precision analysis
This section evaluates the precision of our segmentation, classification, and reconstruction algorithms across different datasets as detailed in Section 2.
4.2.1 ETH Zürich synth3
The ETH Zürich synth3 dataset was selected to showcase the performance of our volumetric method in wall segmentation within a 2.5-dimensional Atlanta-world system. This dataset is frequently used to benchmark the success of interior wall segmentation (Ochmann et al., 2019; Bassier and Vergauwen, 2020; Shi et al., 2019). Although it lacks an exterior scan, this limitation mirrors real-life scenarios where facades are unscanned, thus providing a relevant challenge. Our algorithm successfully classified and generated the geometry for all interior walls, with exterior walls modeled using a predefined thickness. Figure 16 illustrates the resulting IFC model, with the upper ceiling slab hidden to enhance clarity. This model accurately reflects the volumes, positions of openings, and room dimensions. Figure 16 also shows the absolute distance from the points to the wall surfaces, typically within millimeters; a similar accuracy with the deviations up to 2.5 mm in case of walls was achieved also by (Shi et al., 2019). In contrast, the VOX2BIM+ algorithm (Martens and Blankenbach, 2023) exhibited significantly more significant deviations.
Figure 16: IFC model of the ETH Zürich synth3 dataset, automatically generated in Cloud2BIM (left) and a heat map of absolute distances (m) between the point cloud and the resulting IFC model (right).
The dataset also facilitated testing our histogram-based technique complemented by heuristic rules for detecting openings. All internal doors were accurately identified, although the detection was occasionally compromised by furniture or large gaps in point density on the walls. Such errors require manual corrections during model editing in any BIM software.
4.2.2 Kladno station
The Kladno station dataset was used to validate our wall segmentation algorithm on an actual building rather than synthetic data, which was challenging due to the presence of furniture, wall curvatures, and atypical structural elements, such as wall openings with bars. This wall did not conform to standard thickness parameters, highlighting the algorithm’s limitations in processing non-standard elements. The resulting IFC model is depicted in Figure 17. This figure also demonstrates the absolute distance of points from the surfaces, with deviations up to 25 mm. The heat map highlights points over 50 mm away, often associated with staircases, windows, or furnishing elements. These elements were not identified and reproduced, yielding high inaccuracies in the heat map. When detecting zones, one room was incorrectly merged with the other due to failed snapping of imperfectly built walls in a corner of the building, resulting in 24/25 (96%) accuracy (recall Figure 12).
Figure 17: IFC model of the Kladno station dataset, automatically generated in Cloud2BIM (left) and a heat map of absolute distances (m) between the point cloud and the resulting IFC model (right).
4.2.3 Prague hotel
The Prague hotel dataset, a comprehensive scan of a former hotel, showcased our capability to segment extensive building datasets and underscored the accuracy of our ceiling slab segmentation. The model, displayed with transparent walls for clarity, highlights ceiling slabs in yellow (Figure 18). The dataset posed challenges in slab identification due to obstacles and variable point densities, particularly at ground level, where furniture heavily influenced results. Slabs, especially those with only one surface visible, required precise calibration of morphological operations to ensure accurate segmentation. The absolute distances from the point cloud to the slab surfaces, shown in Figure 18, vary up to 25 mm. A few places exhibit the error exceeding 50 mm due to missing point cloud data for the area.
Figure 18: IFC model of the Prague hotel dataset with slabs highlighted in yellow, automatically generated in Cloud2BIM (top) and a heat map of absolute distances (m) between the point cloud and the resulting IFC model (bottom).
4.3 Speed performance analysis
This subsection presents the computational performance results for the test datasets described in Section 2. The evaluations were conducted on a Dell Precision 3591 notebook with an Intel Core Ultra 165H CPU (1.40 GHz), 32 GB RAM, a PM981 NVMe Samsung SSD, and a Windows 11 operating system. The software, developed in Python 3.10, was tested with datasets chosen for their distinctive properties, as detailed in Section 2. Table 8 summarizes the computation times alongside parameters of the resulting models and their point clouds.
Dataset Points Floor area Number of elements Computation time
ETH Zürich synth3 6,905,684 303 m2 74 60 seconds
Kladno station 2,551,721 571 m2 109 30 seconds
Prague hotel 40,152,849 9,348 m2 621 30 minutes
Table 8: Overview of computation results for the test datasets.
While Python’s interpreted nature might not traditionally align with high-performance applications compared to compiled languages like C# or C++, it offered significant advantages in terms of development speed. We leveraged data vectorization capabilities provided by NumPy arrays to boost performance, which proved particularly effective for smaller datasets. For example, the entire reconstruction of the Kladno station, which is relatively modest in scale, was completed in just 32 seconds on standard hardware specifications. In contrast, a more extensive dataset like the Prague hotel, comprising over 40 million points, required 30 minutes to process. This increase in processing time is mainly due to the complexity involved in pairing and joining numerous wall segments. This task could benefit from a compiled language implementation once the software’s development reaches maturity.
Despite its current developmental state, Cloud2BIM outperforms similar software solutions, as evidenced by a comparative analysis with the work by Macher et al. (2017), who did not incorporate opening identification and have not released their code as open-source. Our software demonstrated approximately 7$\times$ the processing efficiency: in a benchmark test, our system processed a 7 million point dataset at a rate of 6.9 million points per minute, compared to the 1.25 million points per minute achieved by the aforementioned study on a 10 million-point dataset. For context, their system required 8 minutes to process a house with approximately 10 million points covering 200 m2 and 19 minutes for the main block of an office building encompassing 1000 m2 with approximately 16 million points (Table 7). Shi et al. Shi et al. (2019) also used the ETH Zürich synth3 datasetand its processing took them 359 seconds without providing the IFC output while we reached the total processing time, including IFC generation, of 60 seconds (Table 8).
4.4 Final remarks and future developments
Cloud2BIM is still evolving, with ongoing enhancements such as the integration of furniture recognition using deep-learning models (Qian et al., 2022). Furthermore, we are developing a graphical user interface that will allow users to smoothly adjust and calibrate input parameters. This interface will feature live visualization, enabling immediate observation of how changes to calibrated parameters affect the model, thus ensuring precise and user-friendly software interaction.
Given the popularity of the RANSAC algorithm for segmenting planar elements from point clouds, one might question why it was not employed in our development. The decision was based on significant concerns regarding performance and the calibration of inputs. While RANSAC is adept at managing noise and outliers, its iterative and random sampling method can inadvertently produce spurious planes that do not correspond to actual planar surfaces (Xu et al., 2016), especially in complex indoor environments. This issue is compounded by the inherent uncertainty in the algorithm’s hypothesis generation phase (Li et al., 2017). Although robust methods are designed to address these inaccuracies, they typically involve point-based operations that are inefficient for indoor settings.
Building on the capabilities of our software, we are exploring technologies to further enrich the BIM models with detailed features and materials not discernible through geometry alone. Inspired by Adán et al. (2018), who utilized RGB data from point clouds to detect minor architectural elements like electrical sockets, we plan to adapt similar image analysis techniques. Although their approach currently lacks IFC export functionality, restricting its integration with BIM tools, it provides a promising avenue for adding realistic details such as color textures and material finishes to our models. This enhancement will not only increase the richness of the visual and functional aspects of the models but also supports comprehensive documentation and precise refurbishment planning.
5 Conclusion
This paper introduces Cloud2BIM, a Python-based open-source automated scan-to-BIM pipeline that converts various point cloud formats into detailed 3D BIM models compliant with the IFC standard. Our approach is distinguished by a volumetric methodology that prioritizes capturing accurately the geometry of building elements over their mere spatial arrangement.
The segmentation and classification process within Cloud2BIM employs specialized algorithms tailored to each building element type, ensuring high accuracy and superior speed compared to other solutions. Basically, Cloud2BIM is able to segment and classify:
  • 1.
    Slabs: The software combines density-based height segmentation, morphological operations, and contour detection to identify floor and ceiling slabs reliably. This method ensures accurate identification even in indoor-only environments without scanning facades or roofs.
  • 2.
    Walls: Wall detection is handled by a hybrid approach that combines the strengths of both room-based and wall-based segmentation. This strategy allows for precise transofmation of point cloud data, accommodating different wall orientations and configurations without the need for exterior scans.
  • 3.
    Openings: Door and window openings are addressed through a histogram-based technique, enhanced with heuristic rules that evaluate the sill positions and the aspect ratio and dimensions of the openings. This approach ensures that each opening is accurately positioned and sized within the BIM model, taking into account architectural standards and construction practices.
  • 4.
    Zones are innovatively defined using actual wall surfaces rather than simple axes. This feature allows for, e.g., detailed definitions of the acoustic, thermal, and usage characteristics of individual rooms. Cloud2BIM creates representation of zones in IFC, opening the possibility for implementing non-graphical information into the BIM model.
Cloud2BIM has been extensively tested on various datasets and has demonstrated high accuracy and superior speed; for example, on the ETH Zürich synth3 dataset, Cloud2BIM achieved millimeter-level accuracy in aligning points to modeled surfaces and surpassed the fastest benchmark solutions by the factor of 7 while providing more functionalities.
As a community-driven platform, Cloud2BIM encourages continuous ehnancement and collaborative development. It aims to integrate future technologies such as machine learning for furniture recognition and augmented reality for enriched contextual data. This integration aims to streamline the creation of detailed BIM documentation and make it more accessible to a wider audience, including small stakeholders and research institutions.
Acknowledgement
We acknowledge the Visualization and MultiMedia Lab at the University of Zurich and Claudio Mura for acquiring the 3D point clouds. We also thank UZH and ETH Zürich for their support in scanning the rooms represented in these datasets. The synth3 dataset they provided was used for validation and testing of our algorithms. Special thanks go to Jan Zeman for consultations and editing and to Rudolf Urban and Zdeněk Svatý for providing us the Kladno station and Prague hotel datasets.
Funding
This work was funded by the European Union under the project ROBOPROX (no. CZ.02.01.01/00/22_008/0004590), by the European Union’s Horizon Europe Framework Programme (call HORIZON-CL4-2021-TWIN-TRANSITION-01-11) under grant agreement No. 101058580, project RECONMATIC (Automated solutions for sustainable and circular construction and demolition waste management), and the Czech Technical University in Prague, grant agreement No. SGS24/003/OHK1/1T/11 (Application of state-of-the-art technologies for enhancing sustainability of the construction sector).